### v1

# Rigidity Criterion for Certain Calabi–Yau families

Abstract.
We prove a new rigidity criterion for families of polarized Calabi-Yau manifolds. Motivated by known non-rigid examples, we conjecture that a family over a quasi-projective curve is rigid if it admits a smooth compactification whose singular fiber has only isolated singularities. We verify this conjecture for singularities with a concentrated mixed Hodge spectrum– class including ordinary double points and cusps. The proof combines an analysis of the vanishing cycle exact sequence and limiting mixed Hodge structure with a tensor-product decomposition of the associated variation of Hodge structures.
2010 Mathematics Subject Classification:
14D22,14C30
1. Introduction
In this paper, we investigate the rigidity of families of Calabi–Yau manifolds over complex algebraic curves. The study of rigidity in algebraic geometry can be traced back to the classical Shafarevich conjecture. While rigidity is well-established for families of hyperbolic curves, the situation for higher-dimensional varieties, particularly Calabi–Yau manifolds, is significantly more subtle. Our primary goal is to provide a criterion for rigidity based on the nature of the singularities developed by the family at the boundary of its compactification.
A rigidity criterion
[…]
The motivation for this work stems from the observation that known examples of nonrigid Calabi–Yau families, such as those constructed by Viehweg and Zuo [VZ05, VZ06], tend to develop singular fibers with singular loci of positive dimension. This leads us to propose that the existence of only isolated singularities in a degenerating family should act as an obstruction to non-rigidity. We formalize this in the following assumption and conjecture:
Assumption 1.1.
Let $f\colon(\mathcal{X},\mathcal{L})\to S$ be a flat family of $n$-dimensional polarized smooth Calabi–Yau manifolds over a quasiprojective curve $S$. Let $\overline{S}$ be a projective smooth compactification of $S$ with a boundary point $0\in\overline{S}\setminus S$. Suppose that the family extends to a flat family $\overline{\mathcal{X}}\to\overline{S}$ such that the total space $\overline{\mathcal{X}}$ is smooth and the singular fiber $X_{0}=f^{-1}(0)$ has only isolated singularities.
Conjecture 1.2.
Under Assumption 1.1, the family $f\colon(\mathcal{X},\mathcal{L})\to S$ is rigid.
Our main result confirms this conjecture for a broad class of isolated singularities:
Theorem 1.3.
In the setting of Assumption 1.1, if $X_{0}$ has only isolated singularities with a concentrated mixed Hodge spectrum, then the family $f\colon\mathcal{X}\to S$ is rigid.
See Definition 2.2 for the precise definition of concentrated mixed Hodge spectrum. As a direct application, since ordinary double points (ODP) and cusp singularities have concentrated spectra (Example 2.3), we obtain:
Corollary 1.4.
In the setting of Assumption 1.1, if $X_{0}$ has only ordinary double points or cusp singularities, then the family $f\colon\mathcal{X}\to S$ is rigid.
Previous works
As aforementioned, for the case of non-isotrivial hyperbolic curves, the family is always rigid. This is part of the geometric version of Shafarevich’s conjecture, proved by Parshin [Par68] and Arakelov [Ara71].
In higher dimensions, the behavior is more complicated due to the existence of nonrigid families. Faltings initially established a rigidity criterion for families of abelian varieties [Fal83], while Peters investigated the problem for abstract variations of Hodge structures (VHS) [Pet90]. Building on these results, Saito-Zucker [SZ91] and Saito [Sai93] classified non-rigid families of K3 surfaces and abelian varieties by analyzing the endomorphism algebras of the corresponding VHS.
These classification results motivated further study of the distribution of non-rigid families in moduli spaces satisfying the Torelli property [CHSZ24]. A corollary is that a non-isotrivial family is rigid if its classifying map covers a very general point in the moduli space. The main theorem of the present paper provides a refinement of this “very general rigidity” specifically for the Calabi–Yau case.
In the Calabi–Yau setting, Liu-Todorov-Yau-Zuo [LTYZ05] established a rigidity criterion based on the non-vanishing of the Yukawa coupling. Subsequently, Zhang [Zha04] proved that a family is rigid if it has a boundary point with maximum unipotent monodromy (i.e. large complex structure limit).
[…]
A conjectural “motivic decomposition”
The known examples of non-rigid families inspire us to ask the following question:
Question 1.5 (Motivic decomposition).
Consider a flat family of $n$-dimensional polarized manifolds over a product surface, $f\colon(\mathcal{X},\mathcal{L})\to S\times B$, whose fibers satisfy the local Torelli property. Here $S$ and $B$ are smooth quasi-projective curves with projective completions $\overline{S}$ and $\overline{B}$. Suppose the family is non-isotrivial along both $S$ and $B$ directions. Assume further that the local system $\mathbb{V}:=R^{n}f_{*}\mathbb{C}_{\mathcal{X}}$ has nontrivial monodromy along the boundary divisor $(S\times(\overline{B}\setminus B))\cup((\overline{S}\setminus S)\times B)$.
Then by Deligne’s theorem, each irreducible factor of $\mathbb{V}$ decompose as tensor product of VHS pulled back from $S$ and $B$ respectively. Are these tensor-product decompositions motivic? More precisely, can one find some product family $(g,h)\colon\mathcal{Y}_{1}\times\mathcal{Y}_{2}\to S^{\prime}\times B^{\prime}$ ($S^{\prime}$ and $B^{\prime}$ are finite coverings of $S$ and $B$) such that after base changing these tensor-products can be realized as irreducible factors of $R^{\cdot}g_{*}\mathbb{C}_{\mathcal{Y}_{1}}\otimes R^{\cdot}h_{*}\mathbb{C}_{\mathcal{Y}_{2}}$?
Recall that in Faltings’ examples of non-rigid families of simple abelian varieties, the base curve is always a compact Shimura curve; consequently, the family has no singular fiber. This explains why the condition requiring non-trivial monodromy along the boundary divisor is necessary in the statement of Question 1.5.
The tensor-product decomposition on VHS imposes restriction on the Mumford-Tate group of $\mathbb{V}$. This gives some Hodge cycle $C$, which, by assuming the Hodge conjecture, corresponds to an algebraic cycle in some self-fiber product of the family $f$. However, it is not clear how $C$ might induce a geometric decomposition of the family.
In another cotext, Viehweg-Zuo use the maximal Higgs field plus the nontrivial local monodromy action along the boundary to decompose the family of abelian varieties into self-product of universal elliptic curves (up to isogeny, cf. [VZ04, Theorem 0.2]). Motivated by these precedents, we intend to investigate Question 1.5 for families with high-weight VHS in the forthcoming work [SYZ].
Under the assumption of Question 1.5, one factor of the predicted product structure in a non-rigid family provides the “room” for the deformation to occur, while the other factor remains fixed over the base $S$. Our criterion essentially posits that the isolated singularities occurring at the boundary are too simple to support the complex, multi-level Hodge structures required by such a non-trivial tensor product decomposition. This observation serves as the primary motivation for Conjecture 1.2.
Methods
The proof proceeds via a contradiction. Suppose the family is nonrigid. By a result of Deligne [Del87], the irreducible factor of the VHS containing the $(n,0)$-part must decompose as a tensor product $\mathbb{V}\cong p_{1}^{*}\mathbb{V}_{1}\otimes p_{2}^{*}\mathbb{V}_{2}$. This decomposition persists for the limiting mixed Hodge structures (MHS) at the boundary point.
On the other hand, we analyze the local behavior near the isolated singularities using the vanishing cycle exact sequence. A key step involves showing that the monodromy operator $T$ acts non-trivially on the $(n,0)$-form $\Omega_{t}$ (Proposition 2.4), thereby establishing the non-triviality of the limiting MHS. The general proof uses deformation theory and the Hodge theory of isolated singularities. For the nodal case specifically, we provide an alternative proof using differential geometric methods from [RT09], [HS17], and [CGPY23]. Notably, the norm bound for the holomorphic volume form due to Collins-Gukov-Picard-Yau yields a better asymptotic estimate for the period integral over vanishing cycles (cf. [CGPY23, Lemma 4.3]). Next we show that the “spreading” effect of the tensor product decomposition (arising from non-rigidity) is incompatible with the "concentration" required by the isolated singularity: Specifically, the tensor product would require multiple non-trivial Hodge filtrations, which contradicts the concentrated mixed Hodge spectrum assumption.
Organization of the Paper
The paper is organized as follows. In Section 2, we review the vanishing cycle sequences and define the concentrated mixed Hodge spectrum for isolated singularities. In Section 3, we analyze the monodromy action on the holomorphic volume form and prove Proposition 2.4 via both Hodge theory and the aforementioned differential geometric approach. Section 4 details the tensor-product decomposition of the limiting MHS and completes the proof of Theorem 1.3. Finally, in Section 5, we discuss the relation of our results to C.-L. Wang’s work on the incompleteness of the Weil–Petersson metric.
Notations and Conventions
Throughout this paper, we work over the field of complex numbers $\mathbb{C}$. A Calabi–Yau manifold is defined as a smooth projective variety with trivial canonical bundle $K_{X}\cong\mathcal{O}_{X}$. For a family $f:\mathcal{X}\to\Delta$, $\mathrm{H}_{\mathrm{lim}}^{n}(X_{t})$ denotes the cohomology of a nearby fiber endowed with the limiting mixed Hodge structure. We denote the semisimple and nilpotent parts of the monodromy $T$ by $T^{ss}$ and $N$, respectively.
Acknowledgment
The first and second authors are grateful to Dingxin Zhang for helpful discussions on the Hodge theory of isolated singularities. The second author is supported by the national key research and development program of China (No. 2022YFA1007100) and NSFC 12201337. The third author is supported by the Key Program (Grant No. 12331002) and the International Collaboration Fund (Grant No. W2441003) of the National Natural Science Foundation of China.
2. Vanishing cycle sequences and singularities
We first study the local behaviour of the family near singular fiber. In this section, we denote by $f\colon\mathcal{X}\to\Delta$ a flat family of $n$-dimensional polarized projective varieties over the unit disk $\Delta=\{t\in\mathbb{C}\mid|t|<1\}$. Assume the family can be completed to a projective base, and $f$ is smooth over $\Delta-\{0\}$. Furthermore, we assume the singular fiber $X_{0}=f^{-1}(0)$ has isolated singularities at $p_{1},\ldots,p_{k}\in X_{0}$. We recall the vanishing cycle exact sequence, see [KL21].
$0\to\mathrm{H}^{n}(X_{0})\xrightarrow{sp}\mathrm{H}_{\mathrm{lim}}^{n}(X_{t})\to\bigoplus_{i=1}^{k}\mathrm{H}^{n}(F_{p_{i}})\to\mathrm{H}^{n+1}(X_{0})\to\mathrm{H}_{\mathrm{lim}}^{n+1}(X_{t})\to 0.$ (1)
Here $F_{p_{i}}$ is the Milnor fiber of the singularity $p_{i}\in X_{0}$. Let $T$ be the monodromy operator arising from $\pi_{1}(\Delta-\{0\})$. Then sequence (1) is $T$-equivariant and respects the mixed Hodge structures. If the total space $\mathcal{X}$ is smooth, then Kerr–Laza [KL21, Corollary 5.7] prove that the morphism $sp$ induces isomorphism
$sp\colon\mathrm{H}^{n}(X_{0})\to\mathrm{H}_{\mathrm{lim}}^{n}(X_{t})^{T}$
where $\mathrm{H}_{\mathrm{lim}}^{n}(X_{t})^{T}$ is the $T$ invariant subspace. Let $T=T^{ss}(I+N)$ be the Jordan decomposition of the monodromy operator, where $T^{ss}$ is the semisimple part and $N$ is the nilpotent part.
In the following, we specialized to the case that $\mathcal{X}\to\Delta$ is family of Calabi–Yau varieties or equivalently $K_{\mathcal{X}}\cong\mathcal{O}_{\mathcal{X}}$. When the singularity is mild with certain rigidity property, we can obtain obstruction to the deformation of the family $\mathcal{X}\to S$. The concentration of Hodge numbers for certain part of $\mathrm{H}_{\mathrm{lim}}^{n}(X_{t})$ determined by $T$ provides the obstruction.
2.1. Mixed spectrum for isolated singularities
First we need the following definition of mixed Hodge spectrum of isolated singularities, see [KL, Definition 1.1].
Definition 2.1.
Let $(X,x)$ be a germ of isolated singularity of dimension $n$. Let $F_{x}$ be the Milnor fiber of the singularity and $T$ be the monodromy operator on $\mathrm{H}^{n}(F_{x})$. Let $T=T^{ss}T^{un}$ be the Jordan decomposition of $T$. Then $\mathrm{H}^{n}(F_{x})$ has eigendecomposition under $T^{ss}$ as $\mathrm{H}^{n}(F_{x})=\bigoplus_{\lambda}\mathrm{H}^{n}_{\lambda}(F_{x})$ with eigenvalues $e^{2\pi\sqrt{-1}\lambda}$ for $\lambda\in[0,1)$. The mixed Hodge strture on $\mathrm{H}^{n}(F_{x})$ gives rise to Hodge–Deligne decompostion on $\mathrm{H}^{n}(F_{x})_{\lambda}=\mathrm{H}^{p,q}_{\lambda}(F_{x})$. The mixed Hodge spectrum of the singularity $(X,x)$ is defined as formal spectrum
$\sigma(X,x)=\sum m_{\alpha,w}[(\alpha,w)]\in\mathbb{Z}[\mathbb{Q}\times\mathbb{Z}]$
where $m_{\alpha,w}=\dim\mathrm{H}^{p,q}_{\lambda}(F_{x})$ with $\alpha=\lambda+p$ and $w=p+q$.
This generalizes to the case of strong smoothings of $X_{0}$ with isolated singularities.
$\sigma(X_{0})=\sum_{i}\sigma(X_{0},p_{i})$
Definition 2.2.
We say the singular fiber $X_{0}$ has concentrated mixed Hodge spectrum if for each $\lambda,w]\in[0,1)\times\mathbb{Z}$, there is at most one $\alpha,w]$ with $\lfloor\alpha\rfloor=\lambda$ such that $m_{\alpha,w}\neq 0$ in the mixed Hodge spectrum $\sigma(X_{0})$.
The mixed spectrum can be calculated explicitly for isolated quasihomogeneous singularities via the combinatorics of the weights, see [Ste77, Ste85, KL].
Example 2.3.
For ordinary double points $(X,x)$ of dimension $n$, the mixed Hodge spectrum is given by
$\sigma(X,x)=\begin{cases}[(\frac{n}{2}+\frac{1}{2},n)],&n\text{ even}\\ [(\frac{n+1}{2},n+1)],&n\text{ odd}\end{cases}$
For cusp singularity $(Y,y)$ of dimension $n$, i.e. locally given by $y_{0}^{2}+y_{1}^{2}+\cdots+y_{n}^{3}=0$, then the mixed Hodge spectrum is given by
$\sigma(Y,y)=\begin{cases}[(\frac{n}{2}+\frac{1}{3},n)]+[(\frac{n}{2}+\frac{2}{3},n)],&n\text{ even}\\ [(\frac{n-1}{2}+\frac{5}{6},n)]+[(\frac{n+1}{2}+\frac{1}{6},n)],&n\text{ odd}\end{cases}$
So if $X_{0}$ has only ordinary double points or cusp singularities as its isolated singularities, then $X_{0}$ has concentrated mixed Hodge spectrum.
2.2. Deformation of top form under strong smoothing
Recall that the family $\mathcal{X}\to\Delta$ viewed as smoothing of $X_{0}$ gives rise to a first order smoothing of each singularity $p_{i}\in X_{0}$. The Kodaira–Spencer class $\theta_{i}\in T^{1}_{X_{0},p_{i}}$ is a strong first-order smoothing because the total space $\mathcal{X}$ is smooth, see [FL24, Definition 0.4]. Under the identification of $T^{1}_{X_{0},p_{i}}$ with $\mathrm{H}^{n}(F_{p_{i}})$ or [FL25, Remark 3.2]. Let $\Omega_{t}$ be the top form on the fiber $X_{t}$. By abuse of notation, $\Omega_{t}$ is also used to denote the class in $\mathrm{H}^{n}(X_{t})$. We claim that the image of $\Omega_{t}$ in $\mathrm{H}^{n}(F_{p_{i}})$ is nonzero for $t\neq 0$ small enough and this also ensures the nontriviality of $T$.
Proposition 2.4.
Let $X_{0}$ be a singular Calabi-Yau variety satisfying Assumption 1.1. Let $f:\,\mathcal{X}\to\Delta$ be a smoothing of $X_{0}$ with $\mathcal{X}$ being smooth. Then for a nearby fiber $X_{t}$ ($t\in\Delta\setminus\{0\}$), the monodromy operator $T$ acts non-trivially on $\mathrm{H}^{n,0}(X_{t})$.
3. Monodromy action and the period integral
The goal of this section is to prove Proposition 2.4. Through this section, we denote by $X_{0}$ the singular Calabi-Yau, $Z:=\mathrm{Sing}(X_{0})$ the singular locus and $U:=X_{0}\setminus Z$ the regular locus.
3.1. Deformation theory
From the local-global spectral sequence of $\mathlcal{Ext}^{i}_{\mathcal{O}_{X_{0}}}$, one receives the following exact sequence
$0\to\mathrm{H}^{1}(T_{X_{0}})\to\mathrm{Ext}^{1}(\Omega^{1}_{X_{0}},\mathcal{O}_{X_{0}})\xrightarrow{\delta}\mathrm{H}^{0}(X_{0},\mathlcal{Ext}^{1}_{\mathcal{O}_{X_{0}}}(\Omega^{1}_{X_{0}},\mathcal{O}_{X_{0}}))\to\mathrm{H}^{2}(T_{X_{0}}),$
where $T_{X_{0}}:=\mathlcal{Hom}_{\mathcal{O}_{X_{0}}}(\Omega^{1}_{X_{0}},\mathcal{O}_{X_{0}})$.
Recall that $\mathrm{Ext}^{1}(\Omega^{1}_{X_{0}},\mathcal{O}_{X_{0}})$ classifies all the first order infinitesimal deformations of $X_{0}$. In fact, we have the following isomorphism:
Lemma 3.1.
$\mathrm{Ext}^{1}(\Omega^{1}_{X_{0}},\mathcal{O}_{X_{0}})\cong\mathrm{H}^{1}(U,T_{U})\cong\mathrm{H}^{1}(U,\Omega^{n-1}_{U})$.
See Proposition 7.6 in [Fri86] or Lemma 4.1 in [FL25] for a proof.
Proposition 3.2.
Let $\xi\in\mathrm{Ext}^{1}(\Omega^{1}_{X_{0}},\mathcal{O}_{X_{0}})$ be the Kodaira-Spencer class associated to a smoothing $\mathcal{X}\to\Delta$ with $\mathcal{X}$ being smooth. Then $\delta(\xi)$ is nonzero in $\mathrm{H}^{0}(X_{0},\mathlcal{Ext}^{1}_{\mathcal{O}_{X_{0}}}(\Omega^{1}_{X_{0}},\mathcal{O}_{X_{0}}))$.
Proof.
By standard argument (cf. Lemma 1.9 in [FL25] for instance), the total space $\mathcal{X}$ being smooth is equivalent to the vanishing of $\mathlcal{Ext}^{1}_{\mathcal{O}_{X_{0}}}(\Omega^{1}_{\mathcal{X}}|_{X_{0}},\mathcal{O}_{X_{0}})$.
Denote by $\mathcal{I}:=\mathcal{O}_{\mathcal{X}}(-X_{0})$ the ideal sheaf of $X_{0}$ in $\mathcal{X}$. Then we have the usual conormal exact sequence $\mathcal{I}/\mathcal{I}^{2}\to\Omega^{1}_{\mathcal{X}}|_{X_{0}}\to\Omega^{1}_{X_{0}}\to 0$. In our case, the conormal sheaf $\mathcal{I}/\mathcal{I}^{2}\cong\mathcal{O}_{X_{0}}$. Hence
$0\to\mathcal{I}/\mathcal{I}^{2}\cong\mathcal{O}_{X_{0}}\to\Omega^{1}_{\mathcal{X}}|_{X_{0}}\to\Omega^{1}_{X_{0}}\to 0.$ (2)
Applying $\mathlcal{Ext}^{i}_{\mathcal{O}_{X_{0}}}$ to (2), one receives the long exact sequence
$0\to T_{X_{0}}\to\mathlcal{Hom}_{\mathcal{O}_{X_{0}}}(\Omega^{1}_{\mathcal{X}}|_{X_{0}},\mathcal{O}_{X_{0}})\to\mathlcal{Hom}_{\mathcal{O}_{X_{0}}}(\mathcal{I}/\mathcal{I}^{2},\mathcal{O}_{X_{0}})\to\mathlcal{Ext}^{1}_{\mathcal{O}_{X_{0}}}(\Omega^{1}_{X_{0}},\mathcal{O}_{X_{0}})\to\mathlcal{Ext}^{1}_{\mathcal{O}_{X_{0}}}(\Omega^{1}_{\mathcal{X}}|_{X_{0}},\mathcal{O}_{X_{0}})=0.$
We obtain the following short exact sequence
$0\to T_{X_{0}}\to\mathlcal{Hom}_{\mathcal{O}_{X_{0}}}(\Omega^{1}_{\mathcal{X}}|_{X_{0}},\mathcal{O}_{X_{0}})\to Q\to 0,$ (3)
where $Q:=\mathrm{Ker}\left(\mathlcal{Hom}_{\mathcal{O}_{X_{0}}}(\mathcal{I}/\mathcal{I}^{2},\mathcal{O}_{X_{0}})\to\mathlcal{Ext}^{1}_{\mathcal{O}_{X_{0}}}(\Omega^{1}_{\mathcal{O}_{X_{0}}},\mathcal{O}_{X_{0}})\right)$. Hence $Q$ is a subsheaf of $\mathlcal{Hom}_{\mathcal{O}_{X_{0}}}(\mathcal{I}/\mathcal{I}^{2},\mathcal{O}_{X_{0}})\cong\mathcal{O}_{X_{0}}$. Since $X_{0}$ is singular, we know that $\mathlcal{Ext}^{1}_{\mathcal{O}_{X_{0}}}(\Omega^{1}_{\mathcal{O}_{X_{0}}},\mathcal{O}_{X_{0}})\neq 0$ and thus $Q\subsetneqq\mathcal{O}_{X_{0}}$. Hence $\mathrm{H}^{0}(Q)=0$. The short exact sequence (3) induces the following
$0\to\mathrm{H}^{1}(T_{X_{0}})\to\mathrm{H}^{1}(\mathlcal{Hom}_{\mathcal{O}_{X_{0}}}(\Omega^{1}_{\mathcal{X}}|_{X_{0}},\mathcal{O}_{X_{0}}))\to\mathrm{H}^{1}(Q).$
Now we use the local-global spectral sequence to obtain the following commutative diagram
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{1}(T_{X_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{1}(\mathlcal{Hom}_{\mathcal{O}_{X_{0}}}(\Omega^{1}_{\mathcal{X}}|_{X_{0}},\mathcal{O}_{X_{0}}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\mathrm{Hom}(\mathcal{I}/\mathcal{I}^{2},\mathcal{O}_{X_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Ext}^{1}(\Omega^{1}_{X_{0}},\mathcal{O}_{X_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi}$$\scriptstyle{\delta}$${\mathrm{Ext}^{1}(\Omega^{1}_{\mathcal{X}}|_{X_{0}},\mathcal{O}_{X_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{0}(\mathlcal{Ext}^{1}_{\mathcal{O}_{X_{0}}}(\Omega^{1}_{\mathcal{O}_{X_{0}}},\mathcal{O}_{X_{0}}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{0}(\mathlcal{Ext}^{1}_{\mathcal{O}_{X_{0}}}(\Omega^{1}_{\mathcal{X}}|_{X_{0}},\mathcal{O}_{X_{0}}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
Note that $\mathrm{Hom}(\mathcal{I}/\mathcal{I}^{2},\mathcal{O}_{X_{0}})$ classifies the first order infinitesimal deformation of $X_{0}$ as a closed subscheme of $\mathcal{X}$. Since $f:\,\mathcal{X}\to\Delta$ is a smoothing of $X_{0}$, its Kodaira-Spencer class $\xi\in\mathrm{Ext}^{1}(\Omega^{1}_{X_{0}},\mathcal{O}_{X_{0}})$ is the image of some nonzero $\widetilde{\xi}\in\mathrm{Hom}(\mathcal{I}/\mathcal{I}^{2},\mathcal{O}_{X_{0}})$. That implies $\psi(\xi)=0$ in $\mathrm{Ext}^{1}(\Omega^{1}_{\mathcal{X}}|_{X_{0}},\mathcal{O}_{X_{0}})\cong\mathrm{H}^{1}(\mathlcal{Hom}_{\mathcal{O}_{X_{0}}}(\Omega^{1}_{\mathcal{X}}|_{X_{0}},\mathcal{O}_{X_{0}}))$. Hence $\xi\not\in\mathrm{H}^{1}(T_{X_{0}})$, which confirms that $\delta(\xi)\neq 0$. ∎
3.2. Proof of Proposition 2.4 by Hodge theory
Recall that $Z=\{p_{1},\dots,p_{k}\}$ is the set of singular points of $X_{0}$. Consider the nearby fiber $X_{t}$ of the smoothing $\mathcal{X}\to\Delta$ of $X_{0}$. Denote by $F_{t}:=\bigcup^{k}_{i=1}F_{p_{i},t}$ the union of Milnor fibers. Let $U_{t}:=X_{t}\setminus F_{t}$ be the complement. Then we have the diffeomorphism $U_{t}\xrightarrow{\simeq}U$, which relates the regular locus of $X_{0}$ and the nearby fiber. Let $L_{t}:=\partial F_{t}$ be the boundary. Then $L_{t}$ is diffeomorphic to $L_{0}$, the union of links of singularities $p_{i}$ (cf. Lemma 6.1 in [Mil68]).
We consider the Mayer–Vietoris sequences
$\cdots\to\mathrm{H}^{n-1}(L_{t})\to\mathrm{H}^{n}(X_{t})\to\mathrm{H}^{n}(U_{t})\oplus\mathrm{H}^{n}(F_{t})\to\mathrm{H}^{n}(L_{t})\to\cdots$ (4)
and
$\cdots\to\mathrm{H}^{n-1}(L_{0})\to\mathrm{H}^{n}(X_{0})\to\mathrm{H}^{n}(U)\to\mathrm{H}^{n}(L_{0})\to\cdots.$ (5)
Lemma 3.3.
The exact sequence (5) is an exact sequence of mixed Hodge structures.
Proof.
Let $\widetilde{X_{0}}\to X_{0}$ be the resolution of singularity of $X_{0}$ and denote $E_{i}$ as the exceptional divisors over $p_{i}$. Then we have the following exact sequence of mixed Hodge structures (MHS):
$\cdots\to\bigoplus_{i=1}^{k}\mathrm{H}^{n-1}(E_{i})\to\mathrm{H}^{n}(X_{0})\to\mathrm{H}^{n}(\widetilde{X_{0}})\to\bigoplus_{i=1}^{k}\mathrm{H}^{n}(E_{i})\to\mathrm{H}^{n+1}(X_{0})\to\cdots.$
Write $E:=\bigcup^{k}_{i=1}E_{i}$. Let $\hat{X_{0}}$ be the germ of $\widetilde{X_{0}}$ along $E$. By [PS08, §6.2], we have the following exact sequence of mixed Hodge structures
[…]
Similarly, one has
[…]
${\mathrm{H}^{k-1}_{E}(\widetilde{X_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\mathrm{H}^{k-1}_{E}(\hat{X_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{k}_{E}(\widetilde{X_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\mathrm{H}^{k}_{E}(\hat{X_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{k-1}(\widetilde{X_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{k-1}(E)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{k}(X_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{k}(\widetilde{X_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{k}(E)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{k-1}(U)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{k-1}(L_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{k}(X_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{k}(U)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{k}(L_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{k}_{E}(\widetilde{X_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\mathrm{H}^{k}_{E}(\hat{X_{0}})}$${\mathrm{H}^{k+1}_{E}(\widetilde{X_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\mathrm{H}^{k+1}_{E}(\hat{X_{0}})}$ (14)
whose columns and the second row are all exact sequences of MHS. This implies (5), as the third row of (14), is also an exact sequence of MHS. ∎
Now we combine (4), (5) with the vanishing cycle exact sequence (1) and obtain the following commutative diagram
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{n-1}(L_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\mathrm{H}^{n}(X_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{sp}$${\mathrm{H}^{n}(U)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{n}(L_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\mathrm{H}^{n-1}(L_{t})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{n}_{\mathrm{lim}}(X_{t})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{n}(U_{t})\oplus\mathrm{H}^{n}(F_{t})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{n}(L_{t})}$${\mathrm{H}^{n}(F_{t})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{n}(F_{t})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{n+1}(X_{0})}$${0}$ (25)
Here $\mathrm{H}^{n}_{\mathrm{lim}}(X_{t})$ denotes the cohomology group $\mathrm{H}^{n}(X_{t})$ endowed with the limiting mixed Hodge structure. By Lemma 3.3 and property of (1), we conclude from (25) that:
Corollary 3.4.
The exact sequence (4) is an exact sequence of MHS.
Now we relate (25) with the deformation theory of $X_{0}$. By Lemma 3.1 and Proposition 3.2, one can enlarge (25) as following
${\mathrm{H}^{1}(T_{X_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\mathrm{Ext}^{1}(\Omega^{1}_{X_{0}},\mathcal{O}_{X_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$$\scriptstyle{\delta}$${\mathrm{H}^{0}(X_{0},\mathlcal{Ext}^{1}_{\mathcal{O}_{X_{0}}}(\Omega^{1}_{X_{0}},\mathcal{O}_{X_{0}}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\mathrm{H}^{2}(T_{X_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\mathrm{H}^{1}_{Z}(\Omega^{n-1}_{X_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{1}(\Omega^{n-1}_{X_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha}$${\mathrm{H}^{1}(U,\Omega^{n-1}_{U})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\beta}$${\mathrm{H}^{2}_{Z}(\Omega^{n-1}_{X_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\gamma}$${\mathrm{H}^{2}(\Omega^{n-1}_{X_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{n-1}(L_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\mathrm{H}^{n}(X_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{sp}$${\mathrm{H}^{n}(U)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{n}(L_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\mathrm{H}^{n+1}(X_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{n-1}(L_{t})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{n}_{\mathrm{lim}}(X_{t})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{n}(U_{t})\oplus\mathrm{H}^{n}(F_{t})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{n}(L_{t})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{n+1}_{\mathrm{lim}}(X_{t})}$ (34)
Proof of Proposition 2.4.
We argue by contradiction. Suppose that the monodromy action of $T$ on $\mathrm{H}^{n,0}(X_{t})$ is trivial.
Let $\Omega_{t}\in\mathrm{H}^{n,0}(X_{t})$ be the holomorphic volume form such that $(\Omega_{t})_{t\in\Delta}$ generates the Hodge line bundle $f_{*}\omega_{\mathcal{X}/\Delta}$. Consider the composed map
$\theta:\,\mathrm{H}^{n,0}(X_{t})\otimes\frac{\partial}{\partial t}\to\mathrm{H}^{n,0}(X_{t})\otimes\mathrm{H}^{1}(T_{t})\xrightarrow{\cup}\mathrm{H}^{n-1,1}(X_{t})=\mathrm{H}^{1}(\Omega^{n-1}_{X_{t}}).$
Then the Kodaira-Spencer class $\theta(\Omega_{t})\in\mathrm{H}^{1}(\Omega^{n-1}_{X_{t}})\subset\mathrm{H}^{n}(X_{t})$ is also $T$-invariant. By the definition of limiting MHS, we know that
$\theta_{0}:=\lim\limits_{t\to 0}\theta(\Omega_{t})\in\mathrm{Gr}^{F_{\cdot}}_{n-1}\mathrm{H}^{n}_{\mathrm{lim}}(X_{t}).$
[…]
Denote by $\xi_{0}$ the image of of $\widetilde{\theta_{0}}$ under the map $\mathrm{H}^{n}(X_{0})\to\mathrm{H}^{n}(U)$. By Lemma 3.3, $\xi_{0}$ is contained in $\mathrm{Gr}^{F_{\cdot}}_{n-1}\mathrm{H}^{n}(U)=\mathrm{H}^{1}(\widetilde{X_{0}},\Omega^{n-1}_{\widetilde{X_{0}}}(\log E))$.
Note that $\Omega^{n-1}_{U}=\Omega^{n-1}_{\widetilde{X_{0}}}(\log E)|_{U}$. Hence one gets the relative cohomology exact sequence
$\cdots\to\mathrm{H}^{1}_{E}(\Omega^{n-1}_{\widetilde{X_{0}}}(\log E))\to\mathrm{H}^{1}(\widetilde{X_{0}},\Omega^{n-1}_{\widetilde{X_{0}}}(\log E))\to\mathrm{H}^{1}(U,\Omega^{n-1}_{U})\to\mathrm{H}^{2}_{E}(\Omega^{n-1}_{\widetilde{X_{0}}}(\log E))\to\cdots,$
which gives the factorization
${\mathrm{H}^{1}(\widetilde{X_{0}},\Omega^{n-1}_{\widetilde{X_{0}}}(\log E))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{n}(U)}$${\mathrm{H}^{1}(U,\Omega^{n-1}_{U})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
Therefore, $\xi_{0}$ can be lifted to an element in $\mathrm{H}^{1}(U,\Omega^{n-1}_{U})\cong\mathrm{Ext}^{1}(\Omega^{1}_{X_{0}},\mathcal{O}_{X_{0}})$, which is exactly the Kodaira-Spencer class associated to the smoothing $f:\,\mathcal{X}\to\Delta$.
[…]
${\mathrm{H}^{1}(\Omega^{n-1}_{\widetilde{X_{0}}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{1}(\Omega^{n-1}_{\widetilde{X_{0}}}(\log E))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{res}}$${\bigoplus^{k}_{i=1}\mathrm{H}^{1}(E_{i},\Omega^{n-2}_{E_{i}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi}$${\mathrm{H}^{n}(\widetilde{X_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{n}(U)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{H}^{n+1}_{E}(\widetilde{X_{0}})\cong\mathrm{H}^{n}(L_{0}),}$ (39)
and the map $\varphi$ factors through $\gamma:\,\mathrm{H}^{2}_{Z}(\Omega^{n-1}_{X_{0}})\to\mathrm{H}^{n}(L_{0})$.
[…]
3.3. Period integral approach to ODP
When the isolated singularities of $X_{0}$ are ordinary double points, there is a differential geometric approach by period integrals. The argument here is from the introduction part of [RT09], see [Wan97, Corollary 3.5] for a proof via Clemens–Schmid exact sequence. Here the nontriviality of the vanishing cycles is from the estimates in geometric analysis of [CGPY23, Lemma 4.3].
Proposition 3.5.
Let $\mathcal{X}$ be a family of Calabi–Yau varieties satisfying the Assumption 1.1 and $X_{0}$ has only ordinary double points as its singularities, then
  1. (1)
    When $n$ is even, then $N=0$, $T^{ss}\neq Id$ and the (1)(-1)-eigenspace $(\mathrm{H}_{\mathrm{lim}}^{n}(X_{t}))^{T=-Id}\neq 0$ has pure Hodge structure with Hodge numbers $h^{p,q}\neq 0$ if and only if $(p,q)=(\frac{n}{2},\frac{n}{2})$.
  2. (2)
    When $n$ is odd, then $N\neq 0$, $T^{ss}=Id$ and $\mathrm{Im}(N)=W_{n-1}(\mathrm{H}_{\mathrm{lim}}^{n}(X_{t}))\neq 0$ has pure Hodge structure with Hodge numbers $h^{p,q}\neq 0$ if and only if $(p,q)=(\frac{n-1}{2},\frac{n-1}{2})$.
Proof of Proposition 3.5.
According to the mixed Hodge structure on vanishing cohomology by Steenbrink [Ste85], the vanishing cohomology $\mathrm{H}^{n}(F_{p_{i}})$ has pure Hodge structure of weight $n$ when $n$ is even and $n+1$ when $n$ is odd.
When $n$ is even, then T=IdT=-Id on $\mathrm{H}^{n}(F_{p_{i}})$ by monodromy of Milnor fibrations. So $T$ is semisimple on $\mathrm{H}_{\mathrm{lim}}^{n}(X_{t})$ by the vanishing cycle sequence (1). Lemma 3.6 ensures the image of $\Omega_{t}$ in $\mathrm{H}^{n}(F_{p})$ is nonzero. So $T$ acts nontrivially on $\mathrm{H}_{\mathrm{lim}}^{n}(X_{t})$ and the (1)(-1)-eigenspace $(\mathrm{H}_{\mathrm{lim}}^{n}(X_{t}))^{T=-Id}\subset\bigoplus_{i=1}^{k}\mathrm{H}^{n}(F_{p})$. So it has pure Hodge structure with Hodge numbers $h^{p,q}\neq 0$ if and only if $(p,q)=(\frac{n}{2},\frac{n}{2})$.
When $n$ is odd, then $T=Id$ on $\mathrm{H}^{n}(F_{p_{i}})$ by monodromy of Milnor fibrations. So $T^{ss}=Id$. Lemma 3.6 ensures the image of $\Omega_{t}$ in $\mathrm{H}^{n}(F_{p})$ is nonzero. On the other hand, the vanishing cohomology $\mathrm{H}^{n}(F_{p})$ has pure Hodge structure of weight $n+1$. So $N^{2}=0$ and $\mathrm{H}_{\mathrm{lim}}^{n}(X_{t})$ has nontrival weights at n1,n,n+1n-1,n,n+1. The weight filtration of $\mathrm{H}_{\mathrm{lim}}^{n}(X_{t})$ is given by $W_{n+1}(\mathrm{H}_{\mathrm{lim}}^{n}(X_{t}))=\mathrm{H}_{\mathrm{lim}}^{n}(X_{t})$, $W_{n}(\mathrm{H}_{\mathrm{lim}}^{n}(X_{t}))=\mathrm{H}^{n}(X_{0})=\ker N$ and $W_{n-1}=\mathrm{Im}(N)$. The operator $N$ induces an isomorphism
$N\colon Gr^{W}_{n+1}(\mathrm{H}_{\mathrm{lim}}^{n}(X_{t}))\xrightarrow{\sim}Gr^{W}_{n-1}(\mathrm{H}_{\mathrm{lim}}^{n}(X_{t})).$
So the conclusion follows. ∎
Now we look at corresponding mixed Hodge structures on $\mathrm{H}^{n}(X_{0})$. Let $\widetilde{X_{0}}\to X_{0}$ be the resolution of singularity of $X_{0}$ such that the exceptional divisors over $p_{i}$ are all n1n-1 dimensional smooth quadrics $E_{i}$. Then we have the following exact sequence of mixed Hodge structures
$\cdots\to\bigoplus_{i=1}^{k}\mathrm{H}^{n-1}(E_{i})\to\mathrm{H}^{n}(X_{0})\to\mathrm{H}^{n}(\widetilde{X_{0}})\to\bigoplus_{i=1}^{k}\mathrm{H}^{n}(E_{i})\to\mathrm{H}^{n+1}(X_{0})\to\cdots$
From the cohomology of quadrics, we know that $\mathrm{H}^{l}(E_{i})\neq 0$ if and only if $l$ is even and $0\leq l\leq 2n-2$. In this case, the Hodge numbers are concentrated in the middle degree. When $n$ is even, then $\mathrm{H}^{n}(X_{0})$ is pure of weight $n$ because $\mathrm{H}^{n-1}(E_{i})=0$. So this also implies $\mathrm{H}_{\mathrm{lim}}^{n}(X_{t})$ is pure of weight $n$ and $N=0$.
[…]
Lemma 3.6.
In the setting of Proposition 3.5 and let $\Omega_{t}$ be the generator of $\mathrm{H}^{0}(X_{t},K_{X_{t}})$ and $\delta_{i}$ be the vanishing cycle near $p_{i}$, then $\int_{\delta_{i}}\Omega_{t}\neq 0$ for $0<|t|<\epsilon$ with $\epsilon$ small enough. In fact $\int_{\delta_{i}}\Omega_{t}\approx c\cdot t^{\frac{n-1}{2}}$ for some constant $c\neq 0$ when $\Omega_{t}$ is the generator of $f_{*}K_{\mathcal{X}/\Delta}$.
Proof of Lemma 3.6.
In fact, the vanishing cycle $\delta_{i}$ can even be represented by special Lagrangian $n$-spheres for $t$ small enough in this case, see [HS17, Corollary A.2] and [CGPY23, Theorem 1.1]. After multiplying a unit complex number, the holomorphic $n$-form $\Omega_{t}$ restricted to $\delta_{i}$ is the volume of the Calabi–Yau metric on $X_{t}$. So $\int_{\delta_{i}}\Omega_{t}\neq 0$ for $0<|t|<\epsilon$ with $\epsilon$ small enough.
The vanishing order of $\int_{\delta_{i}}\Omega_{t}$ near $t$ is a direct consequence of [CGPY23, Lemma 4.3]. In their paper, it is for $n=3$ and the same argument for the zeroth order estimate can be generalized to arbitrary $n$. Choose a coordinate $(x_{0},\cdots,x_{n})$ on $\mathcal{X}$ such that the local singularity at $p_{i}$ is given by $f(x_{0},\cdots,x_{n})=x_{0}^{2}+\cdots+x_{n}^{2}=t$. Then the model residue form $\Omega_{t}$ can be expressed in terms of the residue
$\Omega_{mod,t}=\text{Res}{dx_{0}\wedge\cdots\wedge dx_{n}\over x_{0}^{2}+\cdots+x_{n}^{2}-t}$
Then the zeroth estimate under the Candelas-de la Ossa metric $g_{co,t}$ implies
$|\Omega_{t}-\Omega_{mod,t}|_{g_{co,t}}=O(t^{\frac{1}{2}})$
The vanishing cycle $\delta_{i}$ can be choosen as be
$\delta_{i}=\{(x_{0},\cdots,x_{n})\mid|x_{0}|^{2}+\cdots+|x_{n}|^{2}=|t|,x_{i}\in\sqrt{t}\cdot\mathbb{R}\}$
Then $\int_{\delta_{i}}\Omega_{mod,t}=c\cdot t^{\frac{n-1}{2}}$ for some constant $c\neq 0$ and the integral of $\Omega_{t}$ follows from the zeroth estimate. ∎
4. Tensor-product decomposition of limiting MHS
To prove Theorem 1.3, we argue by contradiction. Suppose the polarized family $f\colon(\mathcal{X},\mathcal{L})\to S$ is non-rigid. Thus it can be deformed to a family of polarized Calabi-Yau manifolds over a product base $S\times B$, where $B$ is some quasiprojective curve. Moreover, the enlarged family is non-isotrivial on both $S\times\{b\}$ and $\{s\}\times B$ for some general points $s\in S$, $b\in B$.
Denote by the enlarged family as $g\colon(\mathcal{Z},\mathcal{L}_{Z})\to S\times B$. Let $\mathbb{V}$ be the primitive part of the $\mathbb{C}$-VHS $R^{n}g_{*}\mathbb{C}_{\mathcal{Z}}$. Then $\mathbb{V}$ is a polarized $\mathbb{C}$-VHS of Calabi-Yau type over $S\times B$. Let $\mathbb{V}_{0}$ be the irreducible factor of $\mathbb{V}$ containing the deepest rank one Hodge filtration $F^{n}V$ of $\mathbb{V}$. Thus $\mathbb{V}_{0}$ is again a $\mathbb{C}$-VHS of Calabi-Yau type over $S\times B$.
Proposition 4.1 (Tensor-product decomposition).
There exist polarized $\mathbb{C}$-VHS $\mathbb{V}_{1}$ over $S$ and $\mathbb{V}_{2}$ over $B$, and a Hodge isometry
$\mathbb{V}_{0}\simeq p^{*}_{1}\mathbb{V}_{1}\otimes p^{*}_{2}\mathbb{V}_{2}.$ (40)
Remark 4.2.
Proposition 4.1 is essentially due to Deligne [Del87], a detailed proof can be found in [VZ05, Proposition 3.3]. Note that the quasi-projectivity of $S$ and $B$ are crucial here since the semisimplicity of $\mathbb{V}_{i}$ are used in the proof.
Therefor, for general points $s\in\overline{S}$ and $b\in B$, one has the isomorphism
$\mathrm{H}_{\mathrm{lim}}(\mathbb{V}_{0,(s,b)})\simeq\mathrm{H}_{\mathrm{lim}}(\mathbb{V}_{1,s})\otimes\mathrm{H}(\mathbb{V}_{2,b})$ (41)
between mixed Hodge structures.
Proof of Theorem 1.3.
We argue by contradiction. Suppose that the family $f\colon(\mathcal{X},\mathcal{L})\to S$ is non-rigid. Then we have the tensor-product decomposition (41) of limiting mixed Hodge structures. Let $T_{0}$ be the monodromy operator on $\mathrm{H}_{\mathrm{lim}}(\mathbb{V}_{0,(s,b)})$. By Proposition 2.4, $T_{0}\neq\mathrm{Id}$. The image of $\mathrm{H}_{\mathrm{lim}}(\mathbb{V}_{0,(s,b)})$ under the vanishing cycle map
$\mathrm{H}_{\mathrm{lim}}(\mathbb{V}_{0,(s,b)})\to\bigoplus_{i=1}^{k}\mathrm{H}^{n}(F_{p_{i}})$
is nontrivial.
On the other hand, the monodromy operator $T_{0}$ can be written as $T_{1}\otimes Id$ from the Hodge isometry (40). So the nontrivial image of $\mathrm{H}_{\mathrm{lim}}(\mathbb{V}_{0,(s,b)})$ with corresponding to certain Hodge spectrum $\lambda,w]\in[0,1)\times\mathbb{Z}$ can also be written as the tensor product of some nontrivial element in $\mathrm{H}_{\mathrm{lim}}(\mathbb{V}_{1,s})$ with Hodge spectrum $\lambda,w_{1}]$ and $\mathrm{H}(\mathbb{V}_{2,b})$ with Hodge spectrum $0,w_{2}]$. Since the variation of Hodge structures $\mathbb{V}_{2}$ is nontrivial, the Hodge spectrum $0,w_{2}]$ corresponds to strictly more than one nontrivial Hodge filtrations. This contradicts to the assumption that $\bigoplus_{i=1}^{k}\mathrm{H}^{n}(F_{p_{i}})$ has concentrated mixed Hodge spectrum. This finishes the proof. ∎
Remark 4.3.
The key point in the proof of Theorem 1.3 is to use the concentrated mixed Hodge spectrum of vanishing cohomology to rule out the tensor-product decomposition of limiting MHS. This idea can also be applied to cases where the limiting MHS already has concentrated mixed Hodge spectrum. For example, if the family $f\colon(\mathcal{X},\mathcal{L})\to S$ has large complex structure limit (or maximum unipotent monodromy) at certain boundary point of $S$, then the deepest Hodge filtration in limiting MHS is equal to the highest possible weight filtration and it is known to be of Hodge-Tate type. So the family $f\colon(\mathcal{X},\mathcal{L})\to S$ must be rigid in this case. Notice that this result is the rigidity of real deformations, which is weaker than the infinitesimal rigidity proved in [Zha04, Theorem 6.11] for polarized Calabi–Yau families with maximum unipotent monodromy.
5. Relation to C.-L. Wang’s result on Weil–Petersson metric
In this section, we discuss the relation between our main result Theorem 1.3 and C.-L. Wang’s result [Wan97, Proposition 2.3] on incompleteness of Weil–Petersson metric for Calabi–Yau families. We specialize to the case of Calabi–Yau threefold with ordinary double points as singularities. Then C.-L. Wang’s [Wan97, Proposition 2.3] implies that the Weil–Petersson metric is incomplete along any smoothing for the singular fiber $X_{0}$.
On the other hand, if the family $f\colon(\mathcal{X},\mathcal{L})\to S$ is non-rigid, then by Proposition 4.1, one has the tensor-product decomposition of $\mathbb{C}$-VHS
$\mathbb{V}_{0}\simeq p^{*}_{1}\mathbb{V}_{1}\otimes p^{*}_{2}\mathbb{V}_{2}.$
Here $\mathbb{V}_{0}$ is the irreducible factor of the primitive part of $R^{3}f_{*}\mathbb{C}_{\mathcal{X}}$ containing the deepest Hodge filtration. By the classfication of all possible nontrivial tensor product decomposition of weight-3 complex Hodge structures of Calabi–Yau type in [VZ06, Proposition 14], the possible period domains corresponding to $\mathbb{C}$-VHS on $\mathbb{V}_{1}$ and $\mathbb{V}_{2}$ are either the complex ball $\mathrm{SU}(1,n)/\mathrm{S}(\mathrm{U}(1)\times\mathrm{U}(n))$ or the type IV symmetric domain $\mathrm{SO}(2,n)/\mathrm{SO}(2)\times\mathrm{SO}(n)$. In both cases, the Hodge metric on the period domains are known to be complete for boundary points with nontrivial unipotent monodromy.
By Proposition 3.5 or [Wan97, Corollary 3.5], the monodromy operator $T$ on $\mathrm{H}_{\mathrm{lim}}^{3}(X_{t})$ satisfies that $N\neq 0$ and $T^{ss}=Id$. So the boundary point corresponding to the singular fiber $X_{0}$ is at infinite distance with respect to the Hodge metric on the period domain of $\mathbb{V}_{1}$. Since the two period domains corresponding to $\mathbb{V}_{1}$ and $\mathbb{V}_{2}$ are both Hermitian symmetric domainis, the Hodge metric is the same as the Weil–Petersson metric up to a constant multiple. Therefore, the Weil–Petersson metric on $S$ is complete along the smoothing for the singular fiber $X_{0}$. This contradicts to C.-L. Wang’s result [Wan97, Proposition 2.3].
Notice that the metric completeness argument here works only for Calabi–Yau threefolds with ordinary double points as singularities. For more general singularities or higher dimensional Calabi–Yau varieties, the monodromy operator may not have nontrivial unitpotent part, or the decomposition factors of the tensor-product decomposition may not correspond to Hermitian symmetric domains. So the metric completeness argument may not introduce the desired contradiction. But more detailed study of the curvature behavior of Weil–Petersson metric near singular fibers may lead to alternative proof of Theorem 1.3 other than the mixed Hodge structure approach.